# MetaSel: A Test Selection Approach for Fine-tuned DNN Models

### v2

@@ Abstract
Deep Neural Networks (DNNs) face challenges during deployment due to data distribution shifts. Fine-tuning adapts pre-trained models to new contexts requiring smaller labeled sets. However, testing fine-tuned models under constrained labeling budgets remains a critical challenge. This paper introduces MetaSel, a new approach, tailored for fine-tuned DNN models, to select tests from unlabeled inputs. MetaSel assumes that fine-tuned and pre-trained models share related data distributions and exhibit similar behaviors for many inputs. However, their behaviors diverge within the input subspace where fine-tuning alters decision boundaries, making those inputs more prone to misclassification. Unlike general approaches that rely solely on the DNN model and its input set, MetaSel leverages information from both the fine-tuned and pre-trained models and their behavioral differences to estimate misclassification probability for unlabeled test inputs, enabling more effective test selection. Our extensive empirical evaluation, comparing MetaSel against 10 state-of-the-art approaches and involving 68 fine-tuned models across weak, medium, and strong distribution shifts, demonstrates that MetaSel consistently delivers significant improvements in Test Relative Coverage (TRC) over existing baselines, particularly under highly constrained labeling budgets. MetaSel shows average TRC improvements of 28.46% to 56.18% over the most frequent second-best baselines while maintaining a high TRC median and low variability. Our results confirm MetaSel’s practicality, robustness, and cost-effectiveness for test selection in the context of fine-tuned models.
Index Terms:
Test Selection, Deep Neural Network, Fine-tuning.
@@ I Introduction
Deep Neural Networks (DNNs) face challenges to widespread adoption, particularly due to data distribution shifts between development and deployment contexts. Training DNNs for new deployment contexts from scratch typically requires extensive labeled input sets, which can be prohibitively expensive. Transfer learning techniques, such as fine-tuning, have emerged as an effective and widely used solution, by leveraging a model trained on a large and diverse input set, fine-tuning it to related but distinct contexts with substantially smaller labeled input sets. Testing fine-tuned DNN models faces similar labeling challenges, requiring effective strategies for selecting test subsets under highly constrained labeling budgets to ensure the reliability and performance of the fine-tuned model while addressing practical resource limitations. In this paper, we introduce MetaSel, an effective test selection approach specifically designed for fine-tuned, classification DNN models that consistently outperforms State-of-the-Art (SOTA), general test selection approaches in terms of misclassification detection given the same budget, particularly under highly constrained budgets.
Several test selection approaches have been introduced for DNN models in recent years [1]. Some of these approaches, such as DeepGD [2], are designed to maximize the identification of misclassification inputs within a predefined labeling budget. Another strategy involves leveraging prioritization approaches to rank inputs based on their probability of being misclassification and thus enable testers to select inputs according to available labeling budgets. An early contribution in this area is the Surprise Adequacy (SA) criteria introduced by Kim et al. [3, 4], which measures how surprising test inputs are to the model compared to its training inputs. Among the most effective and widely used strategies for test input prioritization are uncertainty-based approaches that aim to estimate the model’s confidence in predicting unlabeled test inputs. These strategies can be categorized into probability-based uncertainty metrics, which rely on the model’s output probability vector [5, 6, 1, 7], and neighbor-aware uncertainty metrics, which make use of the input’s nearest neighbors [8, 9]. Recently, learning-based techniques [10, 11] have emerged that entail training an independent model to estimate the probability for an unlabeled test input to be misclassified by DNN Model Under Test (MUT).
@@
To comprehensively evaluate MetaSel, we conducted an extensive empirical study, comparing it against 10 SOTA test input selection baselines. Additionally, we analyzed the impact of varying levels of data distribution shifts on MetaSel’s performance. Specifically, our experiments involved 68 fine-tuned models and corresponding input sets subjected to weak, medium, and strong severity levels of distribution shift between the pre-trained and fine-tuned contexts. Performing this comprehensive study took approximately 39 days, due to both the large number of baselines and subjects, as well as the time-consuming nature of deploying some of the investigated baselines, like the learning-based approach proposed by Demir et al. [10]. Furthermore, we included deeper networks with more complex architectures in our experiments, such as ResNet152, and input sets with a larger number of output classes, such as Cifar-100 with 100 classes. The results demonstrate that MetaSel is a robust solution, consistently outperforming all baselines in selecting more misclassifications within the same budget. Importantly, MetaSel maintains its performance effectiveness across different subjects, selection budgets, and distribution shift levels. Furthermore, our analysis reveals that the second-best performing approach varies across subjects, severity levels, and test selection budgets. This variability highlights MetaSel’s reliability as the only approach that delivers consistently superior performance across diverse conditions.
@@
  • We evaluate MetaSel against 10 SOTA test selection approaches for DNNs as baselines. The results indicate that MetaSel consistently outperforms all baselines by detecting a higher number of misclassified inputs within the same selection budget. Moreover, the findings underscore MetaSel’s reliability, as the second-best performing approach fluctuates across different subjects and selection budgets, further confirming that MetaSel is the only consistent and dependable solution across various scenarios.
@@ II Background
This section provides an overview of four main categories of SOTA prioritization approaches used as baselines in our evaluations. These approaches, known for their effectiveness and efficiency, include surprise-based criteria, probability-based uncertainty scores, neighbor-aware uncertainty scores, and learning-based methods.
For the reasons discussed below, we deliberately exclude some of the existing methods from our evaluations, namely mutation-based [12] and diversity-based [13] techniques, due to their computational impracticality in the context of fine-tuned models and the enormous computational impact they would have on our experiments. PRIMA (PRioritizing test inputs via Intelligent Mutation Analysis) [12], a mutation-based approach proposed by Wang et al. aims to select test inputs that kill the most mutants under the assumption that such inputs are more likely to be mispredicted. However, mutation-based techniques are among the most time-consuming approaches as they require generating mutants and executing each unlabeled test input against all mutants. Adaptive Test Selection (ATS)[13], proposed by Gao et al., adopts an incremental selection strategy to construct a diverse subset of test inputs by iteratively selecting inputs that are the most different from the already selected subset. This approach, however, has been reported as one of the most computationally intensive approaches [8, 14], as it involves calculating pairwise differences between each remaining unlabeled input and the currently selected test subset, thus growing quadratically in cost with the size of the unlabeled test set.
Confirming the above statements, the scalability challenges of PRIMA and ATS have been empirically demonstrated by Hu et al.  [14], who reported a drastic drop in efficiency when these methods were applied to larger datasets such as Cifar-100. Their evaluations show that deploying PRIMA and ATS for test selection on Cifar-100 requires execution times 50 times and 80 times longer, respectively, than Distance-based SA (DSA) [3], one of the most computationally expensive baselines considered in our study. In our specific context, where a large number of unlabeled test inputs can be collected from the target context and is thus expected, as detailed in Section III, these approaches often become impractical. Furthermore, Hu et al. [14] reported that simple probability-based approaches such as DeepGini consistently outperform PRIMA and ATS, which are much more computationally efficient. These findings further reinforce the rationale for excluding these methods from our study.
@@ II-B Probability-based uncertainty metrics
These metrics measure the DNN model’s confidence in predicting a given test input, exclusively relying on the model’s predicted probability distribution [1]. In our experiments, we employ DeepGini [5], Margin [1, 10], and Vanilla-Softmax [6] metrics. Their calculation formulas are presented in Table I. In these formulas, $P_{i}$ is the probability that input $x$ belongs to class $C_{i}$, $C$ is the total number of output classes, and $m$ and $n$ correspond to the indices of the most and second-most probable classes, respectively.
@@ II-C Neighbor-aware uncertainty metrics
Li et al. [8] proposed DATIS (Distance-Aware Test Input Selection), which takes a novel approach to defining the uncertainty score of an unlabeled test input without relying on the DNN model’s output probabilities. Instead, it calculates uncertainty solely based on the ground truth labels of the test input’s nearest neighbors in the labeled training set. To achieve this, DATIS estimates the support of the training data for the DNN’s prediction by computing the normalized sum of the exponentials of the distances between the test input and its nearest training neighbors. The support of the training data for predicting test input ${x}$ as part of class $c\in Y$ based on its $k$-nearest neighbors in the training set is estimated as follows:
@@ II-D Learning-based prioritization
In recent years, learning-based test input prioritization approaches [1] have gained attention for DNN testing. These methods rely on training a model on top of the MUT to estimate the misprediction probability for each test input, leveraging diverse information for enhanced prediction accuracy. Various learning-based approaches have been proposed utilizing different sources of information, such as mutation testing results [12, 16], ensemble model outputs [10], and a test inputs’ nearest neighbors [11]. In our study, we exclude learning-based approaches that either involve high computational costs, such as PRIMA [12] or are not designed for DNNs, such as GraphPrior [16], which was developed specifically for Graph Neural Networks (GNNs). In our evaluations, we consider two SOTA learning-based approaches, namely TestRank [11] and Meta-model [10].
TestRank [11], proposed by Li et al., is one of the well-known learning-based methods that employ a Multi-Layer Perceptron (MLP) to effectively rank test inputs based on their misclassification probability. TestRank involves extracting two categories of features from each test input, namely intrinsic and contextual attributes. Intrinsic attributes are derived from the outputs of the MUT’s logit layer for the given test input while contextual attributes capture the relationships between the test input and its $k$-nearest neighbors. By constructing a graph where nodes represent inputs and edges encode their similarities (measured using a cosine distance metric), TestRank leverages GNNs to extract the contextual attributes of unlabeled test inputs. A simple binary classifier is then trained to predict the probability of an unlabeled input being misclassified by MUT.
@@ III Problem Definition
We also assume that a large target test set $Test_{T}\subset X_{T}$, containing unlabeled inputs, can be collected from the target context. However, labeling all these test inputs is expensive and time-consuming in most practical situations. Therefore, it is essential to prioritize a small-enough subset of unlabeled inputs in $Test_{T}$ to assess the performance of the fine-tuned model while keeping the testing effort within acceptable bounds. Test prioritization approaches assign a score to all inputs in the target test set indicating their relative likelihood of being misclassified by the fine-tuned model, resulting in an ordered list of target test inputs $Test^{ordered}_{T}$. However, ranking the entire target test set is not the primary goal. The ultimate objective in the context of fine-tuned models is, for a given test budget, to select an effective subset of target test inputs $Test^{selected}_{T}$ for labeling and testing the fine-tuned model.
@@ IV MetaSel
In this section, we propose a new learning-based method for ranking and selecting target test inputs to test a fine-tuned model, leveraging its counterpart pre-trained model. Unlike existing general approaches, MetaSel is uniquely tailored for fine-tuned models. While the former offer the advantage of broader applicability across diverse contexts, they fail to leverage the additional artifacts and information available in our specific context, which could significantly enhance the fine-tuned model’s testing.
@@
MetaSel is a DNN model designed to estimate the probability that an unlabeled target test input will be misclassified by the fine-tuned model. We experimented and converged towards a simple DNN model yielding satisfactory results and consisting of a 1D convolutional layer for feature extraction, followed by two or three fully connected layers. In contrast to simpler Neural Networks (NNs) with a shallow single-layer architecture, DNNs have the capability to capture non-linear relationships from complex input features. We define MetaSel as a function $f_{m}:X_{m}\rightarrow[0,1]$ where $X_{m}$ represents the input features and the output $f_{m}(X_{m})$ indicates the predicted probability of an input being misclassified by the fine-tuned model. Consequently, $f_{m}(X_{m})$ values closer to 1 indicate a higher likelihood of misclassification.
To predict the outcome for a given test input, MetaSel relies not only on the output of the fine-tuned model but also on the corresponding output of the pre-trained model for that input, since to some extent knowledge is transferred from the latter to the former. Additionally, to improve MetaSel’s training effectiveness we augment its training set by incorporating a subset of inputs from the source test set $Test_{S}$ that closely aligns with the target data distribution. The detailed procedure for constructing MetaSel, including feature selection and training set augmentation is described in the following sections.

### v4

@@ Abstract
Deep Neural Networks (DNNs) face challenges during deployment due to covariate shift, i.e., data distribution shifts between development and deployment contexts. Fine-tuning adapts pre-trained models to new contexts requiring smaller labeled sets. However, testing fine-tuned models under constrained labeling budgets remains a critical challenge. This paper introduces MetaSel, a new approach tailored for DNN models that have been fine-tuned to address covariate shift, to select tests from unlabeled inputs. MetaSel assumes that fine-tuned and pre-trained models share related data distributions and exhibit similar behaviors for many inputs. However, their behaviors diverge within the input subspace where fine-tuning alters decision boundaries, making those inputs more prone to misclassification. Unlike general approaches that rely solely on the DNN model and its input set, MetaSel leverages information from both the fine-tuned and pre-trained models and their behavioral differences to estimate misclassification probability for unlabeled test inputs, enabling more effective test selection. Our extensive empirical evaluation, comparing MetaSel against 11 state-of-the-art approaches and involving 68 fine-tuned models across weak, medium, and strong distribution shifts, demonstrates that MetaSel consistently delivers significant improvements in Test Relative Coverage (TRC) over existing baselines, particularly under highly constrained labeling budgets. MetaSel shows average TRC improvements of 28.46% to 56.18% over the most frequent second-best baselines while maintaining a high TRC median and low variability. Our results confirm MetaSel’s practicality, robustness, and cost-effectiveness for test selection in the context of fine-tuned models.
@@ I Introduction
Deep Neural Networks (DNNs) face challenges to widespread adoption, particularly due to data distribution shifts between development and deployment contexts. Training DNNs for new deployment contexts from scratch typically requires extensive labeled input sets, which can be prohibitively expensive. Transfer learning techniques, such as fine-tuning, have emerged as an effective and widely used solution by leveraging a model trained on a large and diverse input set, fine-tuning it to related but distinct contexts with substantially smaller labeled input sets. Testing fine-tuned DNN models faces similar labeling challenges, requiring effective strategies for selecting test subsets under highly constrained labeling budgets to ensure the reliability and performance of the fine-tuned model while addressing practical resource limitations. In this paper, we introduce MetaSel, an effective test selection approach specifically designed for fine-tuned, classification DNN models that consistently outperforms State-of-the-Art (SOTA), general test selection approaches in terms of misclassification detection, given the same budget, particularly under highly constrained budgets. MetaSel is specifically designed and evaluated for a scenario where there is a data distribution shift between the development and deployment contexts, while the classification task and thus the model’s output classes remain the same. This scenario, referred to as covariate shift, is framed as a key transfer learning scenario in the comprehensive survey by Pan and Yang [1], who classify covariate shift as a type of common and important domain adaptation.
Several test selection approaches have been introduced for DNN models in recent years [2]. Some of these approaches, such as DeepGD [3], are designed to maximize the identification of misclassification inputs within a predefined labeling budget. Another strategy involves leveraging prioritization approaches to rank inputs based on their probability of being misclassified and thus enable testers to select inputs according to available labeling budgets. An early contribution in this area is the Surprise Adequacy (SA) criteria introduced by Kim et al. [4, 5], which measures how surprising test inputs are to the model compared to its training inputs. Among the most effective and widely used strategies for test input prioritization are uncertainty-based approaches that aim to estimate the model’s confidence in predicting unlabeled test inputs. These strategies can be categorized into probability-based uncertainty metrics, which rely on the model’s output probability vector [6, 7, 2, 8], and neighbor-aware uncertainty metrics, which make use of the input’s nearest neighbors [9, 10]. Recently, learning-based techniques [11, 12] have emerged that entail training an independent model to estimate the probability for an unlabeled test input to be misclassified by DNN Model Under Test (MUT).
@@
To comprehensively evaluate MetaSel, we conducted an extensive empirical study, comparing it against 11 SOTA test input selection baselines. Additionally, we analyzed the impact of varying levels of data distribution shifts on MetaSel’s performance. Specifically, our experiments involved 68 fine-tuned models and corresponding input sets, which were subjected to weak, medium, and strong severity levels of distribution shift between the pre-trained and fine-tuned contexts. Performing such experiments took approximately 48 days (not accounting for interruptions) using a cloud computing environment provided by the Digital Research Alliance of Canada [13]. This extended duration was due to both the large number of baselines and subjects, as well as the time-consuming nature of deploying some of the investigated baselines, like the learning-based approach proposed by Demir et al. [11]. Furthermore, we included deeper networks with more complex architectures in our experiments, such as ResNet152, and input sets with a larger number of output classes, such as Cifar-100 with 100 classes. The results demonstrate that MetaSel is a robust solution, consistently outperforming all baselines in selecting more misclassifications within the same budget. Importantly, MetaSel maintains its performance effectiveness across different subjects, selection budgets, and distribution shift levels. Furthermore, our analysis reveals that the second-best performing approach varies across subjects, severity levels, and test selection budgets. This variability highlights MetaSel’s reliability as the only approach that delivers consistently superior performance across diverse conditions.
@@
  • We evaluate MetaSel against 11 SOTA test selection approaches for DNNs as baselines. The results indicate that MetaSel consistently outperforms all baselines by detecting a higher number of misclassified inputs within the same selection budget. Moreover, the findings underscore MetaSel’s reliability, as the second-best performing approach fluctuates across different subjects and selection budgets, further confirming that MetaSel is the only consistent and dependable solution across various scenarios.
@@
  • Our comprehensive ablation study validates the contribution of each training feature in MetaSel’s performance, thereby confirming the effectiveness of the selected feature set.
@@ II Background
This section provides an overview of the five main categories of SOTA prioritization approaches used as baselines in our evaluations. These approaches, known for their effectiveness and efficiency, include surprise-based criteria, probability-based uncertainty scores, neighbor-aware uncertainty scores, learning-based methods, and diversity-based approaches.
For the reasons discussed below, we deliberately exclude some of the existing methods from our evaluations, namely mutation-based [14] techniques, due to their computational impracticality in the context of fine-tuned models and the enormous computational impact they would have on our experiments. Mutation-based approaches, such as PRIMA (PRioritizing test inputs via Intelligent Mutation Analysis) [14], aim to select test inputs that kill the most mutants under the assumption that such inputs are more likely to be mispredicted. However, mutation-based techniques are among the most time-consuming approaches as they require generating mutants and executing each unlabeled test input against all mutants. The scalability challenges of PRIMA have been empirically demonstrated by Hu et al.  [15], who reported a drastic drop in efficiency when this method was applied to larger datasets such as Cifar-100. Their evaluations show that deploying PRIMA for test selection on Cifar-100 requires execution times 50 times longer, respectively, than Distance-based SA (DSA) [4], one of the computationally expensive baselines considered in our study. In our specific context, where a large number of unlabeled test inputs can be collected from the target context and is thus expected, as detailed in Section III, these approaches often become impractical. Furthermore, Hu et al. [15] reported that simple probability-based approaches such as DeepGini consistently outperform PRIMA, which are much more computationally efficient. These findings further reinforce the rationale for excluding these methods from our study.
@@ II-B Probability-based uncertainty metrics
These metrics measure the DNN model’s confidence in predicting a given test input, exclusively relying on the model’s predicted probability distribution [2]. In our experiments, we employ DeepGini [6], Margin [2], and Vanilla-Softmax [7] metrics. Their calculation formulas are presented in Table I. In these formulas, $P_{i}$ is the probability that input $x$ belongs to class $_{i}, CC$ is the total number of output classes, and $m$ and $n$ correspond to the indices of the most and second-most probable classes, respectively.
@@ II-C Neighbor-aware uncertainty metrics
Li et al. [9] proposed DATIS (Distance-Aware Test Input Selection), which takes a novel approach to defining the uncertainty score of an unlabeled test input without relying on the DNN model’s output probabilities. Instead, it calculates uncertainty solely based on the ground truth labels of the test input’s nearest neighbors in the labeled training set. To achieve this, DATIS estimates the support of the training data for the DNN’s prediction by computing the normalized sum of the exponentials of the distances between the test input and its nearest training neighbors. The support of the training data for predicting test input x{x} as part of class $c\in Y$ based on its $k$-nearest neighbors in the training set is estimated as follows:
@@ II-D Learning-based prioritization
In recent years, learning-based test input prioritization approaches [2] have gained attention for DNN testing. These methods rely on training a model on top of the MUT to estimate the misprediction probability for each test input, leveraging diverse information for enhanced prediction accuracy. Various learning-based approaches have been proposed utilizing different sources of information, such as mutation testing results [14, 18], ensemble model outputs [11], and test inputs’ nearest neighbors [12]. In our study, we exclude learning-based approaches that either involve high computational costs, such as PRIMA [14] or are not designed for DNNs, such as GraphPrior [18], which was developed specifically for Graph Neural Networks (GNNs). In our evaluations, we consider two SOTA learning-based approaches, namely TestRank [12] and Meta-model [11].
TestRank [12], proposed by Li et al., is one of the well-known learning-based methods that employ a Multi-Layer Perceptron (MLP) to effectively rank test inputs based on their misclassification probability. TestRank involves extracting two categories of features from each test input, namely intrinsic and contextual attributes. Intrinsic attributes are derived from the outputs of the MUT’s logit layer for the given test input, while contextual attributes capture the relationships between the test input and its $k$-nearest neighbors. By constructing a graph where nodes represent inputs and edges encode their similarities (measured using a cosine distance metric), TestRank leverages GNNs to extract the contextual attributes of unlabeled test inputs. A simple binary classifier is then trained to predict the probability of an unlabeled input being misclassified by MUT.
@@
II-E Diversity-based prioritization
These approaches aim to select a subset of test inputs that not only has a high potential to detect misclassifications but also exhibits broad input diversity. Robust Test Selection (RTS) introduced by Sun et al. [20], proposes a new probability-tier-matrix-based test selection metric. RTS is a hybrid method that integrates diversity, probability-based uncertainty, and noise filtering into a heuristic test prioritization framework. In this recent approach, the model’s output probability distributions are used to divide test inputs into three distinct sets: noise, suspicious, and successful (correctly classified) inputs. Noise inputs are identified based on their dissimilarity to the model’s training set based on the structural similarity index (SSIM) [21]. To identify successful inputs that are likely to be correctly classified by the model and suspicious inputs that are likely to be misclassified, majority voting from similar training inputs is used. Within each divided set, RTS uses a novel probability-tier-matrix-based metric to measure the output diversity among test inputs and combines it with uncertainty to guide a greedy selection process. Test inputs are then selected sequentially from these sets in a fixed order of priority: first from suspicious, then from successful, and finally from noise inputs.
Adaptive Test Selection (ATS) [17], proposed by Gao et al., adopts an incremental selection strategy to construct a diverse subset of test inputs by iteratively selecting inputs that are the most different from the already selected subset. This approach, however, has been reported as one of the most computationally intensive approaches [9, 15], as it involves calculating pairwise differences between each remaining unlabeled input and the currently selected test subset, thus growing quadratically in cost with the size of the unlabeled test set. Hu et al. [15] reported a significant efficiency drop when applying ATS to larger datasets like Cifar-100, requiring 80 times the execution time of DSA [4], which is already one of the most computationally intensive baselines in our study. More importantly, their findings revealed that simple probability-based methods, such as DeepGini, consistently outperformed ATS. Given that MetaSel consistently outperforms DeepGini across all subjects and test selection budgets, as demonstrated in our evaluations, we exclude ATS from our experiments.
@@ III Problem Definition
We also assume that a large target test set $Test_{T}\subset X_{T}$, containing unlabeled inputs, can be collected from the target context. However, labeling all these test inputs is expensive and time-consuming in most practical situations. Therefore, it is essential to prioritize a small-enough subset of unlabeled inputs in $Test_{T}$ to assess the performance of the fine-tuned model while keeping the testing effort within acceptable bounds. Test prioritization approaches assign a score to all inputs in the target test set, indicating their relative likelihood of being misclassified by the fine-tuned model, resulting in an ordered list of target test inputs $Test^{ordered}_{T}$. However, ranking the entire target test set is not the primary goal. The ultimate objective in the context of fine-tuned models is, for a given test budget, to select an effective subset of target test inputs $Test^{selected}_{T}$ for labeling and testing the fine-tuned model.
@@ IV MetaSel
In this section, we propose a new learning-based method for ranking and selecting target test inputs to test a fine-tuned model, leveraging its counterpart pre-trained model. Unlike existing general approaches, MetaSel is uniquely tailored for fine-tuned models. While the former offers the advantage of broader applicability across diverse contexts, they fail to leverage the additional artifacts and information available in our specific context, which could significantly enhance the fine-tuned model’s testing.
@@
MetaSel is a DNN model designed to estimate the probability that an unlabeled target test input will be misclassified by the fine-tuned model. We experimented and found that a simple DNN architecture consisting of a 1D convolutional layer for feature extraction, followed by two or three fully connected layers, yields consistently strong results while maintaining efficiency. In contrast to simpler Neural Networks (NNs) with a shallow single-layer architecture, DNNs have the capability to capture non-linear relationships from complex input features. To ensure transparency and facilitate reproducibility, MetaSel models constructed for all investigated subjects are publicly available in our replication package [22]. We should note that using a different or more complex architecture may potentially increase MetaSel’s effectiveness while increasing its training and test selection time.
We define MetaSel as a function $f_{m}:X_{m}\rightarrow[0,1]$ where $X_{m}$ represents the input features and the output $f_{m}(X_{m})$ indicates the predicted probability of an input being misclassified by the fine-tuned model. Consequently, $f_{m}(X_{m})$ values closer to 1 indicate a higher likelihood of misclassification.
To predict the outcome for a given test input, MetaSel relies not only on the output of the fine-tuned model but also on the corresponding output of the pre-trained model for that input, since to some extent, knowledge is transferred from the latter to the former. Additionally, to improve MetaSel’s training effectiveness, we augment its training set by incorporating a subset of inputs from the source test set $Test_{S}$ that closely aligns with the target data distribution. The detailed procedure for constructing MetaSel, including feature selection and training set augmentation, is described in the following sections.